Diffusion Morning Report 04/04/12

Ted Berg tberg at stanford.edu
Wed Apr 4 06:43:13 PDT 2012


Good Morning All,
     Looking OK this morning.
     New Issues: FGA2 end cap stuck, No other new issues.
     Leftovers: Thermco Nitride leak, B2H6 regulator change, LTO pull 
and clean. TEOS MFS alarms, SiGe pump out for repair, Thermco TLC 
issues, POCl3 sensor issues,
Epi 2 startup. No other issues at this time.

					
	Date: 	4/4/12 	THICKNESS-Change at 		Priority
Tube Name/Number 	Status 				
Tylan 1 	*up* 	** 	** 	** 	**
Tylan 2 	*up* 	** 	** 	** 	**
Tylan 3 	*up* 	** 	** 	** 	**
Tylan 4 	*up* 	** 	** 	** 	**
Tylan 5-BBr3 	*yellow* 	*tgo sensor issue-ran BBr3 run-observed* 	** 
** 	* 4*
Tylan 6-POCl3 	*yellow* 	*tgo sensor issue tube reinstalled-ran POCl3 
run* 	** 	** 	**
Tylan 7-TEOS2 	*down* 	*mfs error* 	*14.39/30* 	** 	**
Tylan 8-SiGe 	*down* 	*broken tube-bad pump-out for repair* 
*3.4/30* 	** 	**
Tylan 9-Poly 	*up* 	** 	*15.78/30* 	** 	**
Tylan 10-Nitride 	*up* 	** 	*14.22/30* 	** 	**
Tylan 11-LTO 	*down* 	*pull and clean in progress* 	*0./30* 	** 	* 1*
Tylan 12-FGA 	*up* 	** 	** 	** 	**
Thermco 1 	*yellow* 	*no tlc* 	** 	** 	**
Thermco 2 	*yellow* 	*no tlc* 	** 	** 	**
Thermco 3 	*yellow* 	*testing TLC* 	** 	** 	**
Thermco4 	*up* 	** 	** 	** 	**
Thermco Poly1 	*yellow* 	*reg is here installing* 	** 	*
* 	* 2*
Thermco Poly2 	*yellow* 	*reg is here installing* 	** 	** 	**
Thermco Nitride1 	*yellow* 	*leak* 	** 	** 	**
Thermco bpsg 	*up* 	** 	** 	** 	**
AMAT Epi 	*down* 	*starting up,scrubber issues resolved* 	** 	** 	**
TEL Spa 	*yellow* 	*needs test* 	** 	** 	**
Burn Box 	*up* 	** 	** 	** 	**
SEM 4160 	*up* 	** 	** 	** 	**
ZYGO 	*up* 	** 	** 	** 	**
FGA2 	*yellow* 	*end cap stuck* 	** 	** 	* 3*
SiH4 for furnaces 	*on* 	*crossover ok* 	*changeover working well* 	** 	**


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20120404/c5e6e523/attachment.html>


More information about the diffusion mailing list